# Deep Learning for Cross-Domain Data Fusion in Urban Computing: Taxonomy, Advances, and Outlook

### v1

@@ 4.5.3 Diffusion-based Fusion
In recent years, diffusion models Yang et al. (2023a), as an emerging and potent type of deep generative models, have demonstrated state-of-the-art performance across various modalities such as images, speech, and video Cao et al. (2022). Diffusion models can serve as an appropriate framework for data fusion, enabling the seamless incorporation of new modalities due to the existence of conditions Zheng et al. (2023c); Khanna et al. (2023). Simultaneously, the integration of multiple modalities can also enhance the quality of generation.
@@
Urban imagery, as the primary provider of visual information, has borrowed numerous techniques from the general computer vision area and achieved successful applications. DiffusionSat Khanna et al. (2023), inspired by Stable Diffusion Rombach et al. (2022) and ControlNet Zhang et al. (2023d), provided the first large-scale generative foundation model for satellite imagery. It combined data from different modalities including geospatial metadata (latitude, longitude,ground-sampling distance), timestamp, and texts, achieving promising performance on a series of tasks such as super-resolution, temporal generation, and in-painting. The generation of street-view images Gao et al. (2023c); Wang et al. (2024) for spatio-temporal applications, however, remains an underexplored area, which represents a promising research direction in the future.
@@
Aside from common modalities like image, text, and audio, the spatio-temporal modality data covers a broader range of data formats in urban computing, including traffic situations Zhang et al. (2023a), urban flow Zheng et al. (2023c); Xu et al. (2023e); Wen et al. (2023c), and trajectories Yuanshao et al. (2023); Yuan et al. (2022). In this context, the flexibility of diffusion models allows them to potentially exert a more considerable influence.
@@
ChatTraffic Zhang et al. (2023a) creatively introduced fine-grained text in the Text-to-Traffic Generation (TTG) task to adapt to unusual events. Besides, when enhanced with GCN to incorporate the spatial information inherent in the road network, it achieved more accurate and realistic long-term prediction.
@@
DiffSTG Wen et al. (2023c) is the first work that generalizes the diffusion model DDPM to spatio-temporal graphs. As illustrated in Figure 17, it combined the spatio-temporal learning capabilities of STGNNs with the uncertainty measurements of diffusion models. The collection of fine-grained urban flow data is widely recognized as challenging due to the high costs associated with deployment and maintenance, as well as the presence of noise. Diffusion models are suitable tools to generate fine-grained flow maps from the coarse-grained ones Zheng et al. (2023c); Xu et al. (2023e). DiffUFlow Zheng et al. (2023c) is the first generative approach for fine-grained urban flow inference. An ELFetcher module is proposed to utilize various external factors and land features as conditional guidance for the reverse denoising process.
@@
Figure 17: Illustration of DiffSTG and denoising network UGnet, which leverages an Unet-based architecture to capture multi-scale temporal dependencies and the Graph Neural Network (GNN) to model spatial correlations Wen et al. (2023c).
@@
Trajectory data represents another essential type of spatio-temporal data. However, it often raises privacy concerns due to the inclusion of personal geolocation information. One promising solution for this challenge is trajectory generation Yuanshao et al. (2023); Yuan et al. (2022), which aims to generate high-fidelity, privacy-free trajectories. The diffusion model, as a more reliable and robust method of generation than canonical methods, begins to be increasingly explored. DiffTraj Yuanshao et al. (2023) is the first exploration of trajectory generation by the diffusion model, which proposed a Traj-UNet architecture to predict the noise of each diffusion time step. It proved that the step-by-step denoising process of the diffusion model is an appropriate choice due to the stochastic and uncertain characteristics of human activities.
@@ 4.5.4 LLM-enhanced Data Fusion
With the rise of GPTs Brown et al. (2020); Ouyang et al. (2022); Bubeck et al. (2023), the tremendous power of LLMs in a large range of general fields has drawn significant attention, sparking a wave of research paradigm shift and shedding light on artificial general intelligence (AGI). The academia usually terms “large language models (LLM)” for these large-sized PLMs Zhao et al. (2023) due to the scaling laws Kaplan et al. (2020) of Transformer Vaswani et al. (2017) architecture. LLM-enhanced data fusion can be considered as a specific instance of encoder-based alignment, employing LLMs for feature encoding and information interaction. The extensive parameter size of LLMs endows it with potent alignment capabilities. LLMs have multifaceted impacts on the field of urban computing, including work that utilizes LLMs’ geospatial capabilities Roberts et al. (2023); Manvi et al. (2023); Wang et al. (2023d); Xue et al. (2022), applications in remote sensing Hu et al. (2023a); Kuckreja et al. (2023), works in the time series domain Chang et al. (2023); Jin et al. (2023b); Cao et al. (2023); Liu et al. (2023e); Zhou et al. (2023a), etc.
@@
LLMs are able to compress and store geospatial knowledge within the training data. Exploring how to effectively utilize this compressed knowledge and design prompts that stimulate it are areas of current research interest Roberts et al. (2023); Manvi et al. (2023); Wang et al. (2023d); Xue et al. (2022). GPT4GEO Roberts et al. (2023) provided a comprehensive investigation of the extent to which GPT-4 Bubeck et al. (2023) has mastered factual geographic knowledge and its ability to use this knowledge for reasoning. GeoLLM Manvi et al. (2023) proposed that constructing the right prompt is key to extracting geospatial knowledge. Through providing geo-context information near a specific location, LLMs can be fine-tuned to achieve state-of-the-art performance on a variety of large-scale geospatial datasets for tasks such as predicting population density, house price, women’s education, etc. LLM-Mob Wang et al. (2023d) demonstrated the first attempt to apply LLM to modeling human mobility, which reformulated mobility data through historical stays and context stays to introduce long-term and short-term dependencies for prediction and reasoning.
Figure 18: TimeLLM begins by reprogramming the input time series with text prototypes before feeding it into frozen LLM to align the two modalities Jin et al. (2023b).
Large Multi-modal Models (LMMs) Dai et al. ; Liu et al. (2023c); Ye et al. (2023); Zhang et al. (2023e) demonstrate substantial efficacy in information fusion and modal alignment. In the domain of urban computing, the utilization of remote sensing imagery is prevalent for integrating visual elements from a bird’s-eye view perspective.
@@
Inspired by InstructBLIP Dai et al. , RSGPT Hu et al. (2023a) utilized frozen Image Encoder and LLM, while training a lightweight Q-Former Li et al. (2023b) to align the two, achieving state-of-the-art remote sensing image captioning and remote sensing visual question answering downstream tasks. Developed based on LLaVA-v1.5 Liu et al. (2023c), GeoChat Kuckreja et al. (2023) introduced multi-modal instruction-tuning into the remote sensing domain and proposed the first versatile remote sensing Large Vision-Language Model with multitask conversational capabilities.
@@
Following the success of large foundation models in NLP and CV, there are desires in the time series forecasting area to utilize pre-trained LLMs as powerful representation learners Chang et al. (2023); Jin et al. (2023b); Cao et al. (2023); Liu et al. (2023e); Zhou et al. (2023a). LLM4TS Chang et al. (2023) combined patching and channel-independence techniques with temporal encoding, which unlocked the flexibility of pre-trained LLMs without introducing large parameter overhead. As depicted in Figure 18, TimeLLM Jin et al. (2023b) solved the alignment of time series data and natural language modality to unleash the power of LLMs for time series forecasting through Prompt-as-Prefix (PaP), which reprogrammed time series into text prototype representations.
@@ 5.1 Urban Planning
Figure 19: A city exhibits a multi-level structure for urban planning: city-level planning and region-level planning Li et al. (2022b).
@@ 5.1.1 City-level Planning
Multi-sources urban data provides comprehensive information about the operation and variation of cities. Deep learning-based data fusion could utilize the data to sense the city-level variation and understand it as well. This enables urban planners and researchers to understand urban dynamics, such as sensing social events Jiang et al. (2019); Zhao et al. (2022); Guo and Gong (2016), region’s prosperity, city’s vibrancy Liu et al. (2020a) and any important change happened to the city Balsebre et al. (2022). To sense city-level dynamics, Liu et al. (2020a) conducted research for understanding and predicting urban vibrancy evolution based on three data sources: mobile check-in data, POI data, and geographical data. Balsebre et al. (2022) proposed a framework named Geo-ER to match geo-spatial entities from various data sources. Zheng et al. (2023b, a) introduced deep reinforcement learning to autonomously generate road layouts, aiming to connect various locations at the lowest construction expense.
@@
Urban event analysis is crucial for urban planning as variations around a city usually come with urban events. Zhao et al. (2015) proposed a multi-task learning framework to sense and predict urban events from the variation on social media. This framework could effectively train forecasting models for multiple locations simultaneously with shared information by restricting all positions to select a common set of features. Jiang et al. (2019) proposed a recurrent neural network (RNN) based online system named DeepUrbanEvent to understand the crowd dynamic variation for social events. By extracting the deep trend from the current momentary observations on the historical GPS data from citizens, this system could generate an effective prediction for the crowd dynamic trend during a short future time at a big event in a city.
@@
Focusing on the advantages of multi-source data, Zhao et al. (2022) summarized the previous research in event prediction and stressed four challenges for variation understanding and event prediction from multi-source data: 1) geographical hierarchies; 2) hierarchical missing; 3) feature sparsity; 4) difficulty in update with incomplete multiple data. Subsequently, they proposed a multi-source feature learning model based on an Nth-order geo-hierarchy and fused-overlapping group Lasso to handle these challenges. In their methodology, models can be instantly updated from new data from every data source with affordable computational cost, so the urban planner can respond to variation faster with this system.
@@
Road network analysis plays a crucial role in the realm of urban planning, demanding significant attention from professionals in the field. The development of cities leads to continuous variation in the road conditions around a city. It is noteworthy that traditional road maps are unable to cope with these variations. Additionally, conventional surveying methods and database management techniques are inadequate for capturing and managing the rapidly evolving road networks. Yin et al. (2020) pioneered the utilization of multi-source information for the automatic derivation of road attributes. They decided on a multi-task learning framework to extract low-level feature embedding from every data source and applied attention-based fusion to fuse the representations. Based on this work, they could combine the information from GPS trajectory and existing map data and achieve a significant improvement in classification on the OpenStreetMap data in Singapore. Based on the conception of data fusion, Yang et al. (2022) proposed a DuARE system for large-scale automatic extraction by leveraging the GPS trajectory and the satellite images. Driven by its promising performance, DuARE has been deployed in China and has been updating the national road network by 100,000 km every month.
@@ 5.1.2 Region-level Planning
Region detection is the first step for us to understand region-level information. For POI or AOI detection, Xiao et al. (2023) proposed a Contextual Master-Slave Framework (CMSF) that unitizing graph neural network to fusion the POI information and satellite image to detect urban villages. Huang et al. (2023a) fused satellite image, and street-level image with mobility data in their study, and the extra vision data were embedded through a Vision-LSTM network and were demonstrated crucial for region detection through comprehensive ablation study. Their work achieved an overall accuracy of 91.6% in identifying urban villages in Shenzhen, China. Zhao et al. (2016a) carried out a research on integrating information from social media such as Twitter as well as real-world locations. They designed a supervised Bayesian Model (sBM) to analyze the textual information, spatial features, POI information, and user behaviors. Based on this work, researchers are able to find user interests in special regions and understand the regions’ properties.
Region embedding based on multi-source data is the foundation of region-level urban computing. Fu et al. (2019) proposed a multi-view POI network by varying the representation of edges to fusion the geographical distance information and human modality. There are 2 kinds of POI-POI graph networks (shown in Figure 21) in this specialized multi-view framework: 1) distance-based POI-POI networks: the weight of an edge represents the distance between two associated POIs; 2)mobility connectivity-based POI-POI networks: the weight of an edge represents the mobility connectivity between two associated POIs. Zhang et al. (2021b) fused human mobility information with region attributes information with a multi-view joint learning module to infer the land usage of a city and Zhang et al. (2023c) designed a multi-task learning framework called ReMVC to classify regions by land usage and Li et al. (2023d) conducted a research on fusion publicly available building footprint information on OpenStreetMap and POI data for region representation learning. They carried out an experiment in Singapore and New York and demonstrated its efficiency.
Figure 21: An example of the multi-view POI graph networks proposed by Fu et al. (2019).
Bai et al. (2023) also carried out research on geographic representation learning. In their methodology, satellite images and POI information are fused to solve multiple geographic mapping tasks. Besides, Bing et al. (2023) presented a POI embedding method named CatEM, which jointly considered the spatial information and mobility information of POIs and achieved an impressive performance on POI classification tasks. Through efficient region embedding, Wang et al. (2021b, 2023a) proposed a generative framework to generate land-use configuration of a trageted region for urban planning.
@@
Only a few deep learning researches were proposed for region boundary sensing in recent years Wang et al. (2020e); Chen et al. (2022b, 2021b); Vu et al. (2016). The natural boundaries of a city are constantly evolving and adapt to the real needs and functions of the urban area. In fact, accurately sensing these boundaries is crucial for formulating regional policies and adjusting administrative planning to align with the true dynamics of the city. Wang et al. (2020e) conducted a study on mapping the urban boundary of Zhengzhou City, China with its satellite images and POI data and found the result is in great agreement with the boundary ground truth.
Chen et al. (2022b) designed a cross-city federated transfer learning framework named CcFTL that can deal with multi-source data including POIs, road networks, and population density. In their methodology, this framework could not only deal with information from a single city but also transfer within various cities with great robustness. Besides, Chen et al. (2021b) proposed a boundary sensing framework for urban villages that fuse the information from satellite images, mobility from bike-sharing drop-off data, and POIs. This platform was successfully deployed on the government data platform of Xiamen City, China to serve both urban planners and citizens. Some researchers focused on sensing the boundaries of POIs, Vu et al. (2016) proposed a boundary estimation frameworks by pairing the geo-tagged text information on tweets with the real name of POI and analyzing their spatial correlations. This work demonstrated that the spatial distribution of relevant tweets on the platform could reflect the social boundary of the targeted POI.

### v2

@@ 4.5.3 Diffusion-based Fusion
In recent years, diffusion models [325], as an emerging and potent type of deep generative models, have demonstrated state-of-the-art performance across various modalities such as images, speech, and video [30]. Diffusion models can serve as an appropriate framework for data fusion, enabling the seamless incorporation of new modalities due to the existence of conditions [401, 136]. Simultaneously, the integration of multiple modalities can also enhance the quality of generation.
@@
Urban imagery, as the primary provider of visual information, has borrowed numerous techniques from the general computer vision area and achieved successful applications. DiffusionSat [136], inspired by Stable Diffusion [234] and ControlNet [367], provided the first large-scale generative foundation model for satellite imagery. It combined data from different modalities including geospatial metadata (latitude, longitude,ground-sampling distance), timestamp, and texts, achieving promising performance on a series of tasks such as super-resolution, temporal generation, and in-painting. The generation of street-view images [80, 276] for spatio-temporal applications, however, remains an underexplored area, which represents a promising research direction in the future.
@@
Aside from common modalities like image, text, and audio, the spatio-temporal modality data covers a broader range of data formats in urban computing, including traffic situations [358], urban flow [401, 320, 299], and trajectories [356, 354]. In this context, the flexibility of diffusion models allows them to potentially exert a more considerable influence.
@@
ChatTraffic [358] creatively introduced fine-grained text in the Text-to-Traffic Generation (TTG) task to adapt to unusual events. Besides, when enhanced with GCN to incorporate the spatial information inherent in the road network, it achieved more accurate and realistic long-term prediction.
@@
DiffSTG [299] is the first work that generalizes the diffusion model DDPM to spatio-temporal graphs. As illustrated in Figure 18, it combined the spatio-temporal learning capabilities of STGNNs with the uncertainty measurements of diffusion models. The collection of fine-grained urban flow data is widely recognized as challenging due to the high costs associated with deployment and maintenance, as well as the presence of noise. Diffusion models are suitable tools to generate fine-grained flow maps from the coarse-grained ones [401, 320]. DiffUFlow [401] is the first generative approach for fine-grained urban flow inference. An ELFetcher module is proposed to utilize various external factors and land features as conditional guidance for the reverse denoising process.
@@
Figure 18: Illustration of DiffSTG and denoising network UGnet, which leverages an Unet-based architecture to capture multi-scale temporal dependencies and the Graph Neural Network (GNN) to model spatial correlations [299].
@@
Trajectory data represents another essential type of spatio-temporal data. However, it often raises privacy concerns due to the inclusion of personal geolocation information. One promising solution for this challenge is trajectory generation [356, 354], which aims to generate high-fidelity, privacy-free trajectories. The diffusion model, as a more reliable and robust method of generation than canonical methods, begins to be increasingly explored. DiffTraj [356] is the first exploration of trajectory generation by the diffusion model, which proposed a Traj-UNet architecture to predict the noise of each diffusion time step. It proved that the step-by-step denoising process of the diffusion model is an appropriate choice due to the stochastic and uncertain characteristics of human activities.
@@ 4.5.4 LLM-enhanced Data Fusion
With the rise of GPTs [23, 214, 24], the remarkable capabilities of LLMs across a broad spectrum of fields have garnered significant attention, sparking a wave of research paradigm shift and shedding light on artificial general intelligence (AGI). The academia usually terms “large language models (LLM)” for these large-sized PLMs [388] due to the scaling laws [131] of Transformer [268] architecture. LLM-enhanced data fusion can be considered as a specific instance of encoder-based alignment, employing LLMs for feature encoding and information interaction. The extensive parameter size of LLMs endows it with potent alignment capabilities. LLMs have multifaceted impacts on the field of urban computing, including works that utilize LLMs’ geospatial capabilities [232, 195, 291, 322], applications in remote sensing [105, 141], works in the time series domain [34, 126, 29, 181, 402], etc, as illustrated in Figure 19.
Figure 19: Categories of LLM-enhanced data fusion.
@@
LLMs are able to compress and store geospatial knowledge within the training data. Exploring how to effectively utilize this compressed knowledge and design prompts that stimulate it are areas of current research interest [232, 195, 291, 322]. GPT4GEO [232] provided a comprehensive investigation of the extent to which GPT-4 [24] has mastered factual geographic knowledge and its ability to use this knowledge for reasoning. GeoLLM [195] proposed that constructing the right prompt is key to extracting geospatial knowledge. Through providing geo-context information near a specific location, LLMs can be fine-tuned to achieve state-of-the-art performance on a variety of large-scale geospatial datasets for tasks such as predicting population density, house price, women’s education, etc. LLM-Mob [291] demonstrated the first attempt to apply LLM to modeling human mobility, which reformulated mobility data through historical stays and context stays to introduce long-term and short-term dependencies for prediction and reasoning.
@@
Large Multi-modal Models (LMMs) [53, 174, 332, 369] demonstrate substantial efficacy in information fusion and modal alignment. In the domain of urban computing, the utilization of remote sensing imagery is prevalent for integrating visual elements from a bird’s-eye view perspective. Inspired by InstructBLIP [53], RSGPT [105] utilized frozen Image Encoder and LLM, while training a lightweight Q-Former [151] to align the two, achieving state-of-the-art remote sensing image captioning and remote sensing visual question answering downstream tasks. Developed based on LLaVA-v1.5 [174], GeoChat [141] introduced multi-modal instruction-tuning into the remote sensing domain and proposed the first versatile remote sensing Large Vision-Language Model with multitask conversational capabilities.
@@
LLMs have the potential to revolutionize time series analysis [127, 165]. Following the success of large foundation models in NLP and CV, there are desires in the time series forecasting area to utilize pre-trained LLMs as powerful representation learners [34, 126, 29, 181, 402]. LLM4TS [34] combined patching and channel-independence techniques with temporal encoding, which unlocked the flexibility of pre-trained LLMs without introducing large parameter overhead. As depicted in Figure 20, Time-LLM [126] solved the alignment of time series data and natural language modality to unleash the power of LLMs for time series forecasting through Prompt-as-Prefix or Patch-as-Prefix, which reprogrammed time series into text prototype representations.
LLMs approach real-world cross-domain data fusion in the form of agents [127, 310]. In human perception of the world, diverse modalities ultimately converge into language, serving as the medium for expression and communication [239]. Based on essential language capabilities, LLMs can store knowledge and process cross-modal information, thereby continuously receiving feedback and interacting with the environment, demonstrating potential in spatiotemporal domains for comprehensive multimodal data integration. Zhou et al. [404] first explores urban planning through multi-agent collaboration framework, partially demonstrated performance surpassing that of human experts. LLMLight [144] transcends the previous paradigm of using LLMs as assistants to enhance decision-making, by directly employing LLMs as Traffic Signal Control (TSC) agents for decision-making.
Figure 20: Time-LLM begins by reprogramming the input time series with text prototypes before feeding it into frozen LLM to align the two modalities [126].
@@ 5.1 Urban Planning
Figure 21: A city exhibits a multi-level structure for urban planning: city-level planning and region-level planning [154].
@@ 5.1.1 City-level Planning
Multi-sources urban data provides comprehensive information about the operation and variation of cities. Deep learning-based data fusion could utilize the data to sense the city-level variation and understand it as well. This enables urban planners and researchers to understand urban dynamics, such as sensing social events [123, 384, 92], region’s prosperity, city’s vibrancy [177] and any important change happened to the city [13]. To sense city-level dynamics, Liu et al. [177] conducted research for understanding and predicting urban vibrancy evolution based on three data sources: mobile check-in data, POI data, and geographical data. Balsebre et al. [13] proposed a framework named Geo-ER to match geo-spatial entities from various data sources. [397, 395] introduced deep reinforcement learning to autonomously generate road layouts, aiming to connect various locations at the lowest construction expense.
@@
Urban event analysis is crucial for urban planning as variations around a city usually come with urban events. Zhao et al. [385] proposed a multi-task learning framework to sense and predict urban events from the variation on social media. This framework could effectively train forecasting models for multiple locations simultaneously with shared information by restricting all positions to select a common set of features. Jiang et al. [123] proposed a recurrent neural network (RNN) based online system named DeepUrbanEvent to understand the crowd dynamic variation for social events. By extracting the deep trend from the current momentary observations on the historical GPS data from citizens, this system could generate an effective prediction for the crowd dynamic trend during a short future time at a big event in a city.
@@
Focusing on the advantages of multi-source data, Zhao et al. [384] summarized the previous research in event prediction and stressed four challenges for variation understanding and event prediction from multi-source data: 1) geographical hierarchies; 2) hierarchical missing; 3) feature sparsity; 4) difficulty in update with incomplete multiple data. Subsequently, they proposed a multi-source feature learning model based on an Nth-order geo-hierarchy and fused-overlapping group Lasso to handle these challenges. In their methodology, models can be instantly updated from new data from every data source with affordable computational cost, so the urban planner can respond to variation faster with this system.
@@
Road network analysis plays a crucial role in the realm of urban planning, demanding significant attention from professionals in the field. The development of cities leads to continuous variation in the road conditions around a city. It is noteworthy that traditional road maps are unable to cope with these variations. Additionally, conventional surveying methods and database management techniques are inadequate for capturing and managing the rapidly evolving road networks. Yin et al. [339] pioneered the utilization of multi-source information for the automatic derivation of road attributes. They decided on a multi-task learning framework to extract low-level feature embedding from every data source and applied attention-based fusion to fuse the representations. Based on this work, they could combine the information from GPS trajectory and existing map data and achieve a significant improvement in classification on the OpenStreetMap data in Singapore. Based on the conception of data fusion, Yang et al. [324] proposed a DuARE system for large-scale automatic extraction by leveraging the GPS trajectory and the satellite images. Driven by its promising performance, DuARE has been deployed in China and has been updating the national road network by 100,000 km every month.
@@ 5.1.2 Region-level Planning
Region detection is the first step for us to understand region-level information. For POI or AOI detection, Xiao et al. [313] proposed a Contextual Master-Slave Framework (CMSF) that unitizing graph neural network to fusion the POI information and satellite image to detect urban villages. Huang et al. [112] fused satellite image, and street-level image with mobility data in their study, and the extra vision data were embedded through a Vision-LSTM network and were demonstrated crucial for region detection through comprehensive ablation study. Their work achieved an overall accuracy of 91.6% in identifying urban villages in Shenzhen, China. Zhao et al. [381] carried out a research on integrating information from social media such as Twitter as well as real-world locations. They designed a supervised Bayesian Model (sBM) to analyze the textual information, spatial features, POI information, and user behaviors. Based on this work, researchers are able to find user interests in special regions and understand the regions’ properties.
Region embedding based on multi-source data is the foundation of region-level urban computing. Fu et al. [74] proposed a multi-view POI network by varying the representation of edges to fusion the geographical distance information and human modality. There are 2 kinds of POI-POI graph networks (shown in Figure 23) in this specialized multi-view framework: 1) distance-based POI-POI networks: the weight of an edge represents the distance between two associated POIs; 2)mobility connectivity-based POI-POI networks: the weight of an edge represents the mobility connectivity between two associated POIs. Zhang et al. [368] fused human mobility information with region attributes information with a multi-view joint learning module to infer the land usage of a city and Zhang et al. [366] designed a multi-task learning framework called ReMVC to classify regions by land usage and Li et al. [156] conducted a research on fusion publicly available building footprint information on OpenStreetMap and POI data for region representation learning. They carried out an experiment in Singapore and New York and demonstrated its efficiency.
Figure 23: An example of the multi-view POI graph networks proposed by Fu et al. [74].
Bai et al. [10] also carried out research on geographic representation learning. In their methodology, satellite images and POI information are fused to solve multiple geographic mapping tasks. Besides, Bing et al. [20] presented a POI embedding method named CatEM, which jointly considered the spatial information and mobility information of POIs and achieved an impressive performance on POI classification tasks. Through efficient region embedding, Wang et al. [273, 275] proposed a generative framework to generate land-use configuration of a trageted region for urban planning.
@@
Only a few deep learning researches were proposed for region boundary sensing in recent years [294, 37, 40, 270]. The natural boundaries of a city are constantly evolving and adapt to the real needs and functions of the urban area. In fact, accurately sensing these boundaries is crucial for formulating regional policies and adjusting administrative planning to align with the true dynamics of the city. Wang et al. [294] conducted a study on mapping the urban boundary of Zhengzhou City, China with its satellite images and POI data and found the result is in great agreement with the boundary ground truth.
Chen et al. [37] designed a cross-city federated transfer learning framework named CcFTL that can deal with multi-source data including POIs, road networks, and population density. In their methodology, this framework could not only deal with information from a single city but also transfer within various cities with great robustness. Besides, Chen et al. [40] proposed a boundary sensing framework for urban villages that fuse the information from satellite images, mobility from bike-sharing drop-off data, and POIs. This platform was successfully deployed on the government data platform of Xiamen City, China to serve both urban planners and citizens. Some researchers focused on sensing the boundaries of POIs, Vu et al. [270] proposed a boundary estimation frameworks by pairing the geo-tagged text information on tweets with the real name of POI and analyzing their spatial correlations. This work demonstrated that the spatial distribution of relevant tweets on the platform could reflect the social boundary of the targeted POI.